Calculate the frequency of beats produced in air when two sources of sound are activated, one emitting a wavelength of 32 cm and the other of 32.2 cm. The speed of sound in air is 350 m s-1.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Given length of the closed organ pipe, l = 40 cm = 40 × 10^–2 m V base air = 320 So, its frequency p = V/4 l = 320/4 * 40 * 10^-2 = 200 Hertz. As the tuning fork produces 5 beats with the closed pipe, its frequency must be 195 Hz or 205 Hz. Given that, as it is loaded its frequency decreases. So, the frequency of tuning fork = 205 Hz.